# Understanding Video Transformers via Universal Concept Discovery

### v1

@@ 5 Experiments
We evaluate the quality of our concept discovery algorithm quantitatively and qualitatively across a variety of models trained for different tasks.
Datasets. We use two datasets in our experiments: TCOW Kubric [62] and Something-Something-v2 (SSv2) [27]. The former is a synthetic, photo-realistic dataset of 4,000 videos with randomized object location and motion, based on the Kubric synthetic video generator [28]. This dataset is intended for semi-supervised video object segmentation (semi-VOS) through occlusions. SSv2 contains 220,847 real videos intended for finegrained action recognition. Each sample is a crowdsourced video of a person-object interaction (i.e. doing something to something). Unlike many other video classification benchmarks [6, 56], temporal reasoning is fundamental for distinguishing SSv2 actions, making it an ideal choice for analyzing spatiotemporal mechanisms in transformers.
Models. We evaluate four models with public pretrained checkpoints: (i) TCOW [62] trained on Kubric for semi-VOS, (ii) VideoMAE [60] trained on SSv2 for action classification (Supervised VideoMAE), (iii) VideoMAE self-supervised on SSv2 (SSL VideoMAE), and (iv) InternVideo [66], a video-text foundation model trained contrastively on 12M video clips from eight video datasets and 100M image-text pairs from LAION-400M [54]. As TCOW requires a segmentation mask as input, when applying it to SSv2, we hand label the most salient object in the initial frame to use as the query. We focus our analysis on the first two models, and use the last two to validate the universality of our Rosetta concepts.
Implementation details. For all experiments, we run VTCD for 30 randomly sampled videos, and discover concepts from every head at every model layer. Prior work [2] shows Keys produce the most meaningful clusters in self-attention heads, so we focus here on Keys and present results with Queries and Values, together with the rest of the hyperparameters, in the appendix. Our code for the VTCD toolbox allowing to analyze any video transformer representation will be released.
@@
000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.8000.20.20.20.20.40.40.40.4mIoU (%)Positive Perturbation $\downarrow$000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.8Negative Perturbation $\uparrow$
@@
Figure 4: Concept attribution curves for every layer of TCOW trained on Kubric (top) and VideoMAE trained on SSv2 (bottom). We remove concepts ordered from most-to-least (left) or least-to-most important (right). CRIS produces better concept importance rankings than methods based on occlusion (Occ) or gradients (IG).
5.1 Quantitative concept evaluation
To quantitatively confirm the effectiveness of our method, we follow the standard evaluation protocol for concept-based interpretability, and measure the fidelity of the discovered concepts [26, 69, 22]. To this end, we calculate attribution curves [26, 22, 21], where concepts (across all layers) are removed from a model in either most-to-least order (positive perturbation), or least-to-most (negative perturbation). The intuition is that concepts with higher fidelity and more accurate importance scores will have a steeper performance decrease when removing the most important concepts, and vice-versa for the reverse order.
In Figure 4, we plot concept attribution curves of our method for TCOW for Kubric (top) and Supervised VideoMAE for SSv2, targeting 10 randomly sampled classes and averaging results (bottom). In addition, we report results for several baselines: (i) concept removal in a random order, (ii) standard, occlusion-based [21] concept importance estimation, and (iii) a gradient based approach [35, 21]. In all cases, CRIS produces a more viable importance ranking, dramatically outperforming both random ordering and the occlusion baseline. Integrated gradients approach performs similarly to ours for TCOW, but is significantly worse for the action classification VideoMAE model.
Notably, we observe that the performance actually increases for VideoMAE when up to 70% of the least important concepts are removed. Recall that SSv2 is a fine-grained action classification dataset. Our method removes concepts that are irrelevant for the given class, hence increasing the robustness of the model’s predictions. In Section 5.4, we elaborate on this observation and demonstrate how VTCD can be used improve performance and efficiency of any fine-grained video transformer.
Figure 5: The top-3 most important concepts for the TCOW model trained on Kubric (left) and VideoMAE trained on SSv2 for the target class dropping something into something (right). Two videos are shown for each concept and the query object is denoted with a green border in Kubric. For TCOW, the $1^{st}$ and $2^{nd}$ (top-left, middle-left) most important concepts track multiple objects including the target and the distractors. For VideoMAE, the top concept (top-right) captures the object and dropping event (i.e. hand, object and container) while the $2^{nd}$ most important concept (middle-right) captures solely the container. Interestingly, for both models and tasks, the third most important concept (bottom) is a temporally invariant tubelet. See further discussion in Section 5.3 (and video for full results).
5.2 Qualitative analysis
We have seen that the importance assigned to concepts discovered by VTCD aligns well with the accuracy of the model. We now turn to assessing the concepts themselves qualitatively. To this end, in Figure 5 we show two representative videos for the top three most important concepts for the TCOW and VideoMAE models for the class dropping something into something.
For VideoMAE, the most important concepts highlights the object being dropped until the dropping event, at which point both the object and the container are highlighted. The second most important concept clearly captures the container being dropped into, notably not capturing the object itself and making a ring-like shape. These concepts identify an important mechanism of the model that helps it differentiating between similar classes (e.g. dropping something into/behind/in-front of something).
The third most important concept for each model capture similar information, occurring in early layers: a temporally invariant, spatial support. This observation corroborates research [2, 24] suggesting that positional information processing occurs early in the model and acts as a reference frame between the semantic information and the tokens themselves. We now set out to discover concepts that are shared across multiple models, termed Rosetta concepts.
5.3 Rosetta concepts
We begin by applying VTCD to the four models, targeting two classes chosen due to their dynamic nature: rolling something across a flat surface and dropping something behind something. We then mine for Rosetta concepts using the method described in Section 4.1 and setting $\delta=0.15$ and $\epsilon=15\%$ in all experiments. The resulting set of Rosetta 4-concepts contains 40 tuples with an average $R$ score (Equation 7) of 17.1. For comparison, the average $R$ score between all possible 4-concepts is 0.6, indicating the significance of the selected matches. We visualize one of the mined Rosetta 4-concepts in Figure 6, which captures an object tracking representation in the late layers. This result demonstrates that universal representations indeed exist between all four models. We show many more examples of shared concepts on the project web page in video format.
Next, we qualitatively analyze all Rosetta d-concept with $d\in\{2,3,4\}$ at various layers, and show a representative sample in Figure 7. Firstly, we observe that in early layers the models learn spatiotemporal basis representations (Figure 7, left). That is, they decompose the space-time volume of a video into connected regions that facilitate higher-level reasoning in later layers. This is consistent with prior works that showed spatial position is encoded in the early layers of image transformers [2, 24].
Finally, in deeper layers we find concepts that build on top of an object-centric representation to capture specific spatiotemporal events. For example, three out of the four models learn to identify containers before an object has been dropped into them and two models track the object in the hand until it gets dropped. One notable exception here is InternVideo [66], which is primarily trained on static images and has a limited spatiotemporal modeling capacity. On the other hand, we emphasize that these concepts are also found in the self-supervised VideoMAE [60] that was never trained to reason about object-container relationships. This intriguing observation raises the question: can intuitive physics models [7, 67] be learned via large-scale training of video representations?
5.4 Application: Concept pruning with VTCD
Consider a video transformer model trained for $K$ classes, but during deployment, the user is only interested in a subset of classes, $K_{d}\subset K$. Typically, they can either utilize the model as is, with many parameters dedicated to classes outside $K_{d}$, or enhance performance by finetuning the model solely on the subset. Motivated by the findings that different self-attention heads focus on class-specific concepts (Section 5.1), we instead propose a simple method leveraging VTCD to prune unimportant heads to $K_{d}$ and improve model performance and efficiency without finetuning.
@@ 5.4 Application: Concept pruning with VTCD
We choose the VideoMAE model trained on the SSv2 dataset and focus on the six classes where a ‘spill’ takes place (listed in the appendix). We then use CRIS to rank all the heads in VideoMAE according to their effect on performance on those six classes using the training set and report the results for pruning the least important of them on the validation set in Table 1. Pruning the 33% least important heads actually improves the accuracy by 4.3% while reducing FLOPS from 180.5 to 121.5. Further removing 50% of the heads retains the full performance of the original model (+0.7%) and reduces FLOPs to 91.1. These results suggest that one can control the trade-off between performance and computation by choosing the number of heads to remove.
@@ 6 Conclusion
In this work, we introduced VTCD, the first algorithm for concept discovery in video transformers. We experimentally demonstrated that it is capable of extracting human-interpretable concepts from video understanding models and quantifying their importance for the final predictions. Using VTCD, we discovered shared concepts among several models with varying objectives, revealing common processing patterns like a spatiotemporal basis in early layers. In later layers, useful, higher-level representations universally emerge, such as those responsible for object tracking. Large-scale video representation learning is an active area of research at the moment and our approach can serve as a key to unlocking its full potential.
@@
In this appendix, we report additional results, visualizations and implementation details. Note that we include additional video results and corresponding discussions on the project web page. We begin by ablating the tubelet generation component of VTCD and comparing to an existing concept discovery approach in Section 7.1. We then provide statistics of concepts importance distribution between layers in Section 7.2. Next, in Section 7.3, we provide further discussion and qualitative results showing how different concepts are captured in different self-attention heads from the same layer. Finally, we provide further implementation details in Section 8.
@@ 7 Additional results
7.1 Tubelet validation
Recall that, unlike previous methods that partition the inputs into proposals in the pixel space, VTCD generates the proposals via SLIC [1] clustering in the model’s feature space. We ablate this design choice by comparing VTCD with CRAFT [22] - a recent concept discovery approach that uses random cropping for proposal generation, in Table 2.
In particular, we report concept attribution results for the TCOW [62] and VideoMAE [60] models for both our method and CRAFT. In all cases, VTCD result in concepts that are more faithful to the model’s representation. To further isolate the effect of proposals on the performance, we then equip CRAFT with our concept importance estimation approach (shown as ‘CRAFT [22] + CRIS’ in the table). The results confirm our observation from Figure 4 in the main paper that CRIS is superior to the occlusion-based masking used in CRAFT. However, VTCD still outperforms this strong baseline in all settings, validating that generating tubelet proposals in the feature space of the transformer indeed results in concepts that are more faithful to the model’s representation.
@@ 7.2 Quantitative analysis of per-layer concept importance
TCOW VideoMAE
CRAFT [22] 0.174 0.274 0.240 0.300
CRAFT [22] + CRIS 0.166 0.284 0.157 0.607
VTCD (Ours) 0.102 0.288 0.094 0.625
Table 2: Ablation of our tubelet proposal approach via comparison to CRAFT [22] for both TCOW [62] and VideoMAE [60]. Our tubelets result in concepts that are more faithful to the model’s representations even when the baseline is equipped with our concept scoring algorithm (CRAFT [22] + CRIS).
1111222233334444555566667777888899991010101011111111121212120.40.40.40.40.50.50.50.50.60.60.60.60.70.70.70.7LayerImportance (%)Concept Importance
@@ 7.3 Uniqueness of head concepts
Figure 9: Different heads in the same layer capture independent concepts. In layer 5 of TCOW [62], head 6 (top two rows) highlights falling objects, while head 1 (bottom two rows) captures spatial position.
@@ 8 Implementation details
Concept pruning with VTCD. The six classes targeted in the concept pruning application (Section 5.4 in the main paper) are as follows:

### v2

@@ 5 Experiments
We evaluate our concept discovery algorithm quantitatively and qualitatively across a variety of models and tasks.
Datasets. We primarily use two datasets in our experiments: TCOW Kubric [66] and Something-Something-v2 (SSv2) [30]. The former is a synthetic, photo-realistic dataset of 4,000 videos with randomized object location and motion, based on the Kubric synthetic video generator [31]. This dataset is intended for semi-supervised VOS (semi-VOS) through occlusions. SSv2 contains 220,847 real videos intended for fine-grained action recognition. Each sample is a crowdsourced video of a person-object interaction (i.e. doing something to something). Unlike other video classification benchmarks [9, 60], temporal reasoning is fundamental for distinguishing SSv2 actions, making it an ideal choice for analyzing spatiotemporal mechanisms.
Models. We evaluate four models with public pretrained checkpoints: (i) TCOW [66] trained on Kubric for semi-VOS, (ii) VideoMAE [64] trained on SSv2 for action classification (Supervised VideoMAE), (iii) VideoMAE self-supervised on SSv2 (SSL VideoMAE), and (iv) InternVideo [70], a video-text model trained contrastively on 12M clips from eight video datasets and 100M image-text pairs from LAION-400M [58]. As TCOW requires a segmentation as input, when applying it to SSv2, we manually label the most salient object in the initial frame.We focus our analysis on the first two models, and use the last two to validate the universality of our Rosetta concepts.
Implementation details. For all experiments, we run VTCD on 30 randomly sampled videos and discover concepts from all heads and layers. We focus on keys in the self-attention heads, as they produce the most meaningful clusters per prior work [2]. We present results with queries and values on the project page, with further implementation details.
@@
TCOW VideoMAE
Baseline + Occ 0.174 0.274 0.240 0.300
Baseline + CRIS 0.166 0.284 0.157 0.607
VTCD (Ours) 0.102 0.288 0.094 0.625
Table 1: Comparison of our tubelet proposal approach to the widely used random crop tubelets [72, 25, 24] with occlusion-based importance [24] (Baseline + Occ) for both TCOW [66] and VideoMAE [64]. Our tubelets result in concepts that are more faithful to the model’s representations even when the baseline is equipped with our concept scoring algorithm (Baseline + CRIS).
5.1 Quantitative evaluation
This section presents a quantitative validation of VTCD’s key components compared to various baselines. We first assess our tubelet proposal methodology, followed by a comparison of CRIS with other concept importance methods. We follow the standard evaluation protocol, and measure the fidelity of the discovered concepts [29, 74, 25]. To this end, we calculate attribution curves [29, 25, 24], where concepts are removed from a model in either most-to-least order (positive perturbation), or least-to-most (negative perturbation). The intuition is that concepts, and corresponding importance scores, with higher fidelity will have a steeper performance decrease when removing the most important concepts, and vice-versa for the reverse order.
Tubelet validation. Unlike previous methods that partition the inputs into proposals in the pixel space, VTCD generates the proposals via SLIC [1] clustering in the model’s feature space. We ablate this design choice by comparing VTCD with a strong random cropping baseline used by recent image-based concept discovery methods (ConceptSHAP [72], CRAFT [25], Lens [24]), shown as ‘Baseline + Occ’ in Table 1. In all cases, VTCD yields concepts that are more faithful to the model’s representation. To further isolate the effect of proposals on the performance, we equip the baseline with our concept importance estimation approach (shown as ‘Baseline + CRIS’). We observe that VTCD still outperforms this strong baseline in all settings, validating that generating tubelet proposals in the feature space of the transformer indeed yields concepts that are more faithful to the model’s representation.
[…]
@@ 5.1 Quantitative evaluation
Figure 4: Attribution curves for every layer of TCOW trained on Kubric (top) and VideoMAE trained on SSv2 (bottom). We remove concepts from most-to-least (left) or least-to-most important (right). CRIS produces better concept importance than methods based on single concept occlusions (Occ) or gradients (IG).
@@
Figure 5: The top-3 most important concepts for the TCOW model trained on Kubric (left) and VideoMAE trained on SSv2 for the target class dropping something into something (right). Two videos are shown for each concept and the query object is denoted with a green border in Kubric. For TCOW, the $1^{st}$ and $2^{nd}$ (top-left, middle-left) most important concepts track multiple objects including the target and the distractors. For VideoMAE, the top concept (top-right) captures the object and dropping event (i.e. hand, object and container) while the $2^{nd}$ most important concept (middle-right) captures solely the container. Interestingly, for both models and tasks, the third most important concept (bottom) is a temporally invariant tubelet. See Section 5.3 for further discussion (and the video for full results).
Concept important evaluation. In Figure 4, we plot concept attribution curves of our method for TCOW for Kubric (top) and supervised VideoMAE for SSv2, targeting 10 randomly sampled classes and averaging results (bottom). In addition, we report several baselines: (i) concept removal in a random order, (ii) standard, occlusion-based concept importance estimation [24], and (iii) a gradient based approach [37, 24]. In all cases, CRIS produces a more viable importance ranking, dramatically outperforming both random ordering and the occlusion baseline. The integrated gradients method performs similarly to ours for TCOW, but is much worse for the action classification VideoMAE.
@@
Notably, we observe that the performance actually increases for VideoMAE when up to 70% of the least important concepts are removed. Recall that SSv2 is a fine-grained action classification dataset. VTCD removes concepts that are irrelevant for the given class, hence increasing the robustness of the model’s predictions. We further quantify this effect in Table 2, where we focus on the six classes where a ‘spill’ happens (listed in the appendix). We then use CRIS to rank the heads in VideoMAE according to their effect on performance using the training set and report the results for pruning the least important ones on the validation set. Pruning 33% of the heads actually improves the accuracy by 4.3% while reducing FLOPS from 180.5 to 121.5. Further removing 50% of the heads retains the original performance (+0.7%) and reduces FLOPs to 91.1. These results show that one can tune the trade-off between performance and computation using VTCD for pruning.
5.2 Qualitative analysis
We have seen that the importance assigned to concepts discovered by VTCD aligns with the accuracy of the model. We now assess the concepts qualitatively. To this end, Figure 5 shows two representative videos for the top three most important concepts for the TCOW and VideoMAE models for the class dropping something into something.
For VideoMAE, the most important concept highlights the object until it is dropped, after which the object and its container are highlighted. The second concept clearly captures the container, but notably not the object itself, making a ring-like shape. These concepts identify an important mechanism for differentiating similar classes (e.g. dropping something into/behind/in-front of something).
The third concept for each model captures similar information, occurring in early layers: a temporally invariant, spatial support. This corroborates research [2, 27] suggesting that positional information processing occurs early in the model, acting as a reference frame between semantic information and the tokens themselves. We now turn to Rosetta concepts: concepts shared across multiple models.
5.3 Rosetta concepts
We begin by applying VTCD to the four models, targeting two classes chosen due to their dynamic nature: rolling something across a flat surface and dropping something behind something. We then mine Rosetta concepts following Section 4.1, with $\delta=0.15$ and $\epsilon=15\%$ in all experiments. The resulting Rosetta 4-concepts contains 40 tuples with an average $R$-score (Equation 7) of 17.1. Note the average $R$-score between all possible 4-concepts is 0.6, indicating the significance of the selected matches. Figure 6 visualizes one mined Rosetta 4-concept, which captures an object tracking representation in the late layers. This demonstrates that universal representations indeed exist between all four models. Our project website shows more shared concepts as videos.
Next, we qualitatively analyze all Rosetta d-concept with $d\in\{2,3,4\}$ at various layers. Figure 7 shows a representative sample. In early layers, we find that the models learn spatiotemporal basis representations (Figure 7, left): they decompose the video space-time volume into connected regions, facilitating higher-level reasoning in later layers. This is consistent with prior works that show spatial position is encoded in early layers of image transformers [2, 27].
Features VTCD VTCD + SAM [38]
VideoMAE-SSL 45.0 68.1
VideoMAE 43.1 66.6
InternVideo 45.8 68.0
Table 3: We apply VTCD to discover object tracking concepts and evaluate them with mIoU on the DAVIS’16 validation set.
We validate the accuracy of the discovered object-centric concepts on the DAVIS’16 video-object segmentation (VOS) benchmark [52]. To this end, we run VTCD on the training set and select concepts that have the highest overlap with the groundtruth tracks. We then use them to track new objects on the validation set and report results in Table 3 (see appendix for implementation details). In this analysis, we solely evaluate representations that are not directly trained for VOS. While solid performance is achieved by all representations, segmentation accuracy is limited by the low resolution of the concept masks. To mitigate this, we introduce a simple refinement step that samples random points inside the masks to prompt SAM [38]. The resulting approach, shown as ‘VTCD + SAM’, significantly improves performance. We anticipate that future developments in video representation learning will automatically lead to better VOS methods with the help of VTCD.
Finally, deeper layers contain concepts that build on object-centric representations to capture spatiotemporal events. For example, three models learn to identify containers before an object is dropped into them and two models track the object in the hand until it is dropped. One notable exception is InternVideo [70], which is primarily trained on images with limited spatiotemporal modeling. Perhaps surprisingly, these concepts are also found in the self-supervised VideoMAE [64], which was never trained to reason about object-container relationships. This raises the question: Can intuitive physics models [10, 71] be learned via large-scale video representation training?
@@ 6 Conclusion
In this work, we introduced VTCD, the first algorithm for concept discovery in video transformers. We empirically demonstrated that it is capable of extracting human-interpretable concepts from video understanding models and quantifying their importance for the final predictions. Using VTCD, we discovered shared concepts among several models with varying objectives, revealing common processing patterns, such as a spatiotemporal basis in early layers. In later layers, useful, higher-level representations universally emerge, such as those responsible for object tracking. Large-scale video representation learning is an active area of research at the moment [6, 4, 7] and our approach can serve as a key to unlocking its full potential.
@@
In this appendix, we report additional results, visualizations and implementation details. We first conduct validation experiments for VTCD to show that the discovered concepts align with human-interpretable groundtruth labels in Section 7.1. We then provide statistics of concepts importance distribution between layers in Section 7.2. Next, in Section 7.3, we provide further discussion and qualitative results showing how different concepts are captured in different self-attention heads from the same layer. We provide further implementation details in Section 8. Finally, we discuss limitations of VTCD in Section 9. Note that we include additional video results and corresponding discussions on the project web page.
@@ 7 Additional results
7.1 Concept validation
Directly measuring concept accuracy is not possible in an open world approach as we don’t know a priori what concepts should be present in a model. There are, however, special cases where we can directly measure the accuracy of some of the concepts. For example, TCOW is trained to track through occlusions, so we can expect to find concepts that correspond to the target object and containers/occluders in it. We perform this evaluation for VTCD and the random crop baseline [72, 25, 24] and report the mIoU between the best found concepts and the groundtruth masks in Table 4 (top). These results validate the accuracy of VTCD, which is able to reach up to 94% of the performance of the fully-supervised TCOW by discovering concepts in its intermediate representations.
@@ 7.3 Uniqueness of head concepts
Method Target Occluders Containers
Baseline 3.0 31.5 43.3
VTCD (Ours) 19.2 69.7 73.8
TCOW (supervised) 36.8 76.8 78.2
Table 4: Evaluating the accuracy of object tracking concepts found in the TCOW model by VTCD and the random crop baseline used in other recent methods [72, 25, 24].
1111222233334444555566667777888899991010101011111111121212120.40.40.40.40.50.50.50.50.60.60.60.60.70.70.70.7LayerImportance (%)Concept ImportanceTCOW - VOSVideoMAE - AR
@@ 8 Implementation details
Concept pruning with VTCD. The six classes targeted in the concept pruning application (Table 2 in the main paper) are as follows:
@@
Figure 9: Different heads in the same layer encode concepts capturing different phenomena. In layer 5 of TCOW [66], head 6 (top two rows) highlights falling objects, while head 1 (bottom two rows) captures spatial position.
Semi-Supervised VOS with VTCD. To evaluate VOS performance on the DAVIS’16 benchmark [52] with any pretrained video transformer, we first identify self-attention heads that encode object-centric concepts on the training set. Specifically, we calculate the mIoU between every concept found by VTCD (i.e. the set of tubelets belonging to that concept) and the groundtruth labels for each training video. We then record the heads in which the best performing concepts came from. Next, we run VTCD on the validation set using only the heads containing object-centric concepts. To select the final concept for evaluation, we choose the one with the highest mIoU with the first frame label (i.e. the query mask).
To generate tubelets for an entire video, we simply use non-overlapping sliding windows and run SLIC on the temporally concatenated features. We also leverage SAM [38] for post-processing the tubelets generated by VTCD. Note that SAM can take as input a set of points, a bounding box, or both. For each frame and corresponding mask from the VTCD tubelet, we generate the smallest bounding box surrounding the mask. We also calculate the centroid of the mask and then sample two points from a Gaussian centered at the centroid, with covariance $(L/10,W/10)$ where $L$ and $W$ are the length and width of the mask, respectively. We then pass both the box and points to SAM ViT-h to produce the post-processed mask for that frame.
9 Limitations
One limitation of our method is the need to manually set the SLIC compactness hyper-parameter. Additionally, the compactness property of SLIC makes it challenging to capture concepts that are not spatially localized. Finally, there is a high computational requirement for computing the Rosetta score in Equations 7 and 8 as $D$ grows.